OpenStudy (anonymous):

is The following is a geometric sequence. 1, -2, 4, -8, 16,…

OpenStudy (anonymous):

it is times -2

OpenStudy (anonymous):

yeah,common ratio is -2

OpenStudy (anonymous):

You can check it yourself.. Just find the common ratio: Common Ratio: -2/1 = 4/(-2) = (-8)/4 = -2 So this is a Geometric Sequence..